A firm has total cost of TC(y)=y²+1 and marginal cost of MC(y)=2y. What is the firm's profit at price $8?

Group of answer choices

10

0

15

neither one is correct

25
Answer #1

Answer:

neither one is correct
reason:

TC = y^2+1; MC = 2y So, ATC =(y^2+1)/y

the firm produces where ATC = MC.
(y^2+1)/y =  2y

y^2+1 = 2y^2 So, y^2 = 1; y = 1.

Profit = Total revenue - total cost

= (1*8) - (1^2+1)

= 8 - 2 =6 ( no option has this answer)